Question Hep says that less than 80% of adults in a certain country think that Of 200 addS nthat country, 79% think that be vaccinated A1a : 0 05 es there enough evidence to support the researchers claim? Complete parts (a) through (e) below Ho and Ha (Type an integer or a decimal Do not round ) A. Less than % of adults in the country think that healthy children should be required to be vaccinated B. More than 1% of adults in the country think that healthy children should be required to be vaccinated. of adults in the country think that healthy children should be required to be vaccinated ОС O D. % The percentage of adults in the country who think that healthy children should be required o be va crated is not 1% (Round to two decimal places as needed) ов. Нор Ha p> Hop> Ha P E. 204 AM 11/19/2018 39
Answer #1

Here we have n = 200 , \widehat{p} = 0.76

a) Less than 80% of adults in the country think that healthy children should be required to be vaccinated.

b) The hypothesis are :

H0 : P \geq 0.80 v/s H1: P < 0.80

c) The test statistic is

z=\frac{\widehat{p}-P}{\sqrt{\frac{P(1-P)}{n}}}

  =\frac{0.76-0.80}{\sqrt{\frac{0.80(1-0.80)}{200}}}

= -1.41

p value = p ( z < -1.41 ) = 0.0793

( look in to standard normal probability table across -1.4 and 0.01 )

Here p value > \alpha ( 0.05 )

Hence we failed to reject null hypothesis.

Conclusion : There is no enough evidence to support researchers claim.